OpenStudy (anonymous):

Solve the equation by any method. x^(2)+√(5x)-2=0 Which method could I use for this problem?

OpenStudy (kirbykirby):

A good thing to do first would be to get rid of your square root by sending it to the other side of the equality, then square both sides

zepdrix (zepdrix):

Hmm I'm not sure how you would approach this one... http://www.wolframalpha.com/input/?i=x%5E2%2B%285x%29%5E%281%2F2%29-2%3D0 If you go down to where it says "Real Solution:" Click the box that changes it to exact form, you'll see that the answer is crazy -_-

OpenStudy (kirbykirby):

LOL haha. Yeah indeed it appears this one would be tricky as @zepdrix has shown..

OpenStudy (tkhunny):

...unless you just wrote it incorrectly. Did you mean \(\sqrt{5}x\) or did you REALLY mean \(\sqrt{5x}\)? Very different animals.

OpenStudy (anonymous):

Ah! Crap.... I wrote it wrong you where right @tkhunny its supposed to not include x in it. Thank you for pointing that out

OpenStudy (tkhunny):

Perfect. Did you try the Quadratic Formula?

OpenStudy (amoodarya):

Quadratic Formula

OpenStudy (anonymous):

Okay I did that and I don't know. That radical is annoying me. Not sure if this is right but I got \[\frac{ -\sqrt{5}\pm \sqrt{13} }{ 2 }\] ?

OpenStudy (tkhunny):

When you start with irrational coefficients, you should expect that. Only a miracle will clear it out to integers for you.

OpenStudy (anonymous):

Ah I got it right! Yay! Thank you everyone.